OpenStudy (elonasushchik):

Hey! Write the equation of the line that passes through (2, 4) and has a slope of −1 in point-slope form. A.) x − 4 = −1(y − 2) B.) x − 2 = −1(y − 4) C.) y − 2 = −1(x − 4) D.) y − 4 = −1(x − 2)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (elonasushchik):

k then..

pooja195 (pooja195):

Plug in the info

pooja195 (pooja195):

To get your answer

OpenStudy (chosenmatt):

^^^

OpenStudy (elonasushchik):

k I think I got C. not sure

OpenStudy (nincompoop):

cool question do you need a clear picture on how to understand these things?

OpenStudy (nincompoop):

I will help you when pooja is done typing

OpenStudy (elonasushchik):

Thnx luv u pooji:) kisses E>

OpenStudy (nincompoop):

help you help yourself of course

pooja195 (pooja195):

Rember (x,y) <---it is in that order right? so x=2 y=4

pooja195 (pooja195):

Try again :)

OpenStudy (elonasushchik):

huh?

pooja195 (pooja195):

C is incorrect.

OpenStudy (elonasushchik):

Ohh its D then

pooja195 (pooja195):

Good :)

OpenStudy (elonasushchik):

thanx hun

pooja195 (pooja195):

;)

OpenStudy (nincompoop):

and the guessing lesson started

pooja195 (pooja195):

@nincompoop that wasn't a guess. She got confused on the coordinate part so i explained.

OpenStudy (elonasushchik):

yeah..

OpenStudy (nincompoop):

ahuh